# Image Complexity-Aware Adaptive Retrieval for Efficient Vision-Language Models

### v1

@@ Abstract
Vision transformers in vision-language models apply uniform computational effort across all images, expending 175.33 GFLOPs (ViT-L/14) whether analysing a straightforward product photograph or a complex street scene. We propose ICAR (Image Complexity-Aware Retrieval), which enables vision transformers to use less compute for simple images whilst processing complex images through their full network depth. The key challenge is maintaining cross-modal alignment: embeddings from different processing depths must remain compatible for text matching. ICAR solves this through dual-path training that produces compatible embeddings from both reduced-compute and full-compute processing. This maintains compatibility between image representations and text embeddings in the same semantic space, whether an image exits early or processes fully. Unlike existing two-stage approaches that require expensive reranking, ICAR enables direct image-text matching without additional overhead. To determine how much compute to use, we develop ConvNeXt-IC, which treats image complexity assessment as a classification task. By applying modern classifier backbones rather than specialised architectures, ConvNeXt-IC achieves state-of-the-art performance with 0.959 correlation with human judgement (Pearson) and 4.4× speedup. Evaluated on standard benchmarks augmented with real-world web data, ICAR achieves 20% practical speedup while maintaining category-level performance and 95% of instance-level performance, enabling sustainable scaling of vision-language systems.
@@ 1 Introduction
Modern search engines use Vision Transformers (ViT) within vision-language models to match images with text queries, enabling applications from web image search to e-commerce product matching. ViTs process images through multiple layers in sequence, with each layer adding more detailed understanding of the image [raghu2021vision].
@@
Standard ViT systems have a fundamental inefficiency: they use the same amount of compute for all images, regardless of the complexity of the image. For example, a 24-layer ViT-L/14 uses the same 175.33 GFLOPs [ilharco_gabriel_2021_5143773] of compute whether it’s processing a simple product photo or a complex street scene. At web scale, where major platforms process billions of images daily [GoogleIO2024], this uniform processing can lead to a substantial waste of computational resources. The logical solution is adaptive computation: allowing simple images to use less compute while complex images use more compute.
@@
Adaptive computation can be achieved through early exit: given that ViTs process images through a sequence of layers that progressively refine extracted features, simple images can stop processing at intermediate layers [raghu2021vision], while complex images continue through the full network. Adaptive computation for vision-language models faces a fundamental technical challenge: maintaining cross-modal alignment when embeddings are produced from different processing depths. While adaptive computation has been extensively studied for single-modality tasks [teerapittayanon2016branchynet], the adaptive vision-language field remains underexplored. Specifically, no existing work addresses adaptive computation for image-text retrieval tasks.
@@
We address this by proposing ICAR (Image Complexity-Aware Retrieval), which leverages image complexity to determine the processing path for each image (Figure 1). ICAR introduces dual-path training that produces compatible embeddings from both early reduced compute and full processing paths. This enables simple images to use fewer layers whilst complex images process through the full network, all while maintaining cross-modal alignment through unified embedding spaces, eliminating the reranking overhead that existing approaches have. Evaluated on standard benchmarks (Flickr30K [young2014flickr30k]/MS-COCO [lin2014microsoft]) augmented with 100K real-world samples from LAION-5B, ICAR achieves 20% practical speedup while maintaining category-level performance and 95% of instance-level performance.
@@
To enable ICAR’s routing decisions, we develop ConvNeXt-IC (ConvNeXt-Image Complexity). While existing methods treat complexity assessment as a representation learning problem requiring specialised architectures, we instead recognise it as a classification task. Therefore, by using modern classification backbones pretrained on ImageNet [deng2009imagenet] rather than self-supervised representation learning approaches, ConvNeXt-IC achieves state-of-the-art performance (0.959 Pearson correlation, 4.4 speedup). We validate across both established academic benchmarks (IC9600) and real-world web data (LAION), demonstrating that our approach generalises beyond controlled settings. The following are the contributions of the paper:
@@
First adaptive computation method for image-text retrieval: We introduce ICAR, which addresses the cross-modal alignment challenge through dual-path training that produces compatible embeddings from different processing depths, enabling direct text matching without reranking overhead.
@@
Reconceptualising complexity assessment as classification: We demonstrate that image complexity assessment is better suited as a classification task, rather than representation learning. By treating it as classification with modern backbones (ConvNeXt-V2) instead of building specialised complexity assessment architectures, we achieve state-of-the-art performance (0.959 Pearson correlation) with 4.4 speedup over existing representation learning methods.
Differential computational requirements: We observe that category-level and instance-level retrieval have different computational needs: category matching requires only coarse semantic features while instance matching demands fine-grained discrimination, enabling targeted efficiency optimisations with 20% practical speedup.
@@
This paper is organised as follows: Section 2 discusses related work in image complexity assessment and adaptive computation. Section 3 presents our ConvNeXt-IC complexity detection method with validation on established benchmarks and real-world data. Section 4 details the ICAR architecture and dual-path training approach. Section 5 presents experimental results and analysis of retrieval performance and efficiency gains. Finally, Section 6 concludes with key findings and future directions.
@@ 2.1 Image Complexity Assessment
Image complexity assessment determines how difficult an image is for humans to understand or describe [snodgrass1980standardized]. Psychology research identified key factors: object count, spatial relationships, visual diversity, and background richness [olivia2004identifying, forsythe2009visual, purchase2012predicting]. Early computational approaches used hand-crafted features like entropy and compression ratios [stamps2002entropy, machado2015computerized, rosenholtz2007measuring], but these worked only on small datasets.
Current deep learning methods treat complexity assessment as a representation learning problem. ICNet [feng2022ic9600] uses a dual-branch architecture combining detail and context pathways with spatial attention modules. ICCORN [guo2023image] applies ordinal regression to ICNet features, explicitly modelling the ordinal structure of complexity ratings. MICM [li2025unlocking] introduces motion-inspired assessment using image-to-video generation with hierarchical alignment losses. CLICv2 [liu2025clicv2] uses self-supervised pretraining on complexity-related visual patterns.
Related image quality assessment methods explore similar perceptual challenges but focus on quality rather than complexity. HyperIQA [su2020blindly] employs content-adaptive hyper networks for quality assessment, while CLIPIQA [wang2022exploring] leverages vision-language representations from CLIP [radford2021learning]. TOPIQ [chen2024topiq] adopts a top-down framework modelling human visual perception characteristics.
These methods follow a common pattern: start with classification baselines, then add complexity-specific architectures for incremental improvements. However, they test only against outdated networks like AlexNet [krizhevsky2012imagenet] (2012) and ResNet [he2016deep] (2016), leaving a large gap unexplored. This raises an open question whether architectural specialisation provides genuine benefits, or whether stronger classification backbones would close the performance gap without the need for additional components. We address this gap using ConvNeXt-V2 (2023) to test whether modern classification architectures outperform specialised complexity assessment designs.
@@ 2.2 Adaptive Computation in Machine Learning
Without adaptive computation, a vision-language model such as OpenCLIP’s ViT-L/14 uses 175.33 GFLOPs for each sample, regardless of whether the sample is challenging or easy to represent [ilharco_gabriel_2021_5143773]. The vision encoder consumes 162.03 GFLOPs while text processing requires only 13.3 GFLOPs, yet both apply the same computational effort to all inputs. This uniform processing creates inefficiencies where simple product photos receive the same 24-layer transformer processing as complex street scenes with dozens of objects [dosovitskiy2020image].
Single-modality methods have explored adaptive computation to address this inefficiency by adjusting processing based on input difficulty. BranchyNet [teerapittayanon2016branchynet] introduced early exits with classifiers at intermediate layers, enabling confident predictions to exit early while difficult cases continue through all layers. Several methods adapt early exits for Vision Transformers through patch processing, token removal, and routing strategies [yin2022avit, chen2023cf, rao2021dynamicvit, bolya2023token, riquelme2021scaling].
@@
Limited work exists on adaptive computation for vision-language models. DeeCap [fei2022deecap] and MuE [tang2023you] target generative tasks, while HiVLP [chen2022hivlp] uses hierarchical features for retrieval but requires re-ranking due to separate embedding spaces. No existing work enables direct single-stage retrieval with embedding compatibility across processing depths.
At web scale, where platforms process billions of images daily, efficiency gains have immediate infrastructure impact through reduced energy costs and faster response times. No existing work addresses adaptive computation for image-text retrieval, leaving this critical efficiency challenge unsolved.
@@ 3 Image Complexity Detection Method
This section presents and validates our proposed ConvNeXt-IC image complexity detection method. This method acts as a component for our proposed image-text retrieval method ICAR, the integration of which is discussed in Section 4.
@@ 3.1 Proposed Method: ConvNeXt-IC
To adapt image-text retrieval models to use more or less compute based on image complexity, we need automatic complexity assessment. Real-world web data exhibits a wide range of complexity levels. Figure 3 illustrates this with representative cases from the LAION-5B dataset [schuhmann2022laion].
These examples illustrate the range of complexity that methods must handle. Existing approaches treat complexity assessment as a representation learning problem, building specialised architectures to capture complexity patterns [feng2022ic9600, guo2023image, li2025unlocking]. However, the core challenge is discriminating between different levels of visual complexity. We therefore treat complexity assessment as a classification problem, using efficient classification backbones instead of specialised complexity architectures.
We therefore fine-tune ConvNeXt-V2-N (224x224) [woo2023convnext], the smallest of the ConvNeXt-V2 ImageNet-22K [deng2009imagenet] family of foundation classification models. This choice prioritises efficiency for real-time routing decisions while maintaining the discriminative power needed for accurate complexity assessment. We implement ConvNeXt-IC with two configurations: (1) As a regression model (using a single linear layer predicting continuous scores); and (2) as a binary classification model (using a lightweight two-layer MLP with cross-entropy loss). Both configurations are fine-tuned end-to-end using standard data augmentation (random crops, horizontal flips, colour jittering) on the benchmark datasets.
3.2 Validation on IC9600 Benchmark
To validate our approach, we evaluate ConvNeXt-IC on IC9600 [feng2022ic9600] in its regression head setting, the established benchmark for image complexity assessment. IC9600 contains 9,600 images labelled with complexity scores from 1 (simple) to 5 (complex). We measure performance using standard regression metrics: Pearson Correlation Coefficient (PCC) and Spearman Rank Correlation Coefficient (SRCC) assess how well predicted scores correlate with human ratings, while Root Mean Square Error (RMSE) measures prediction accuracy. We also report inference speed (images/second) to evaluate computational efficiency. Table 1 shows our results compared to existing methods.
@@ 3.2 Validation on IC9600 Benchmark
| HyperIQA | 0.935 | 0.926 | 0.204 | 61 |
| CLIPIQA | 0.898 | 0.897 | 0.078 | 83 |
| TOPIQ | 0.944 | 0.938 | 0.049 | 125 |
| CLICv2 | 0.933 | 0.927 | - | - |
| ICNet | 0.949 | 0.945 | 0.053 | 397 |
| MICM | 0.953 | 0.943 | 0.049 | 0.01 |
| ICCORN | 0.955 | 0.951 | - | - |
| ConvNeXt-IC (Ours) | 0.959 | 0.956 | 0.045 | 1744 |
ConvNeXt-IC achieves state-of-the-art performance: 0.959 PCC and 0.956 SRCC, outperforming the previous best method ICCORN (0.955 PCC, 0.951 SRCC) while delivering 1,744 images/second throughput. Most notably, ConvNeXt-IC surpasses methods that were specifically architecturally designed for image complexity detection, like ICNet (0.949 PCC) and MICM (0.953 PCC). This demonstrates that strong, correctly chosen backbones which are then fine-tuned can outperform specialised architectures, providing another instance of the current trend in machine learning where data and compute together can outperform sophisticated architectural design [halevy2009unreasonable, zhai2022scaling].
3.3 Domain Shift Spot Check on LAION Subset
While IC9600 provides a controlled validation benchmark, its images are sourced from curated academic datasets. To verify that ConvNeXt-IC generalises to the noisy, real-world web content that ICAR is targeting, we conduct a domain shift spot check on LAION data.
We sample 3,000 images from LAION-5B [schuhmann2022laion] and manually annotate them into complexity classes of ‘complex’ and ‘simple’. The annotation process naturally resulted in the classes being distributed as 49.5% simple and 50.5% complex. Note that, this annotation process was conducted solely by the lead researcher of this paper. Therefore, while the primary validation of ConvNeXt-IC was conducted on IC9600 as discussed in the previous subsection, experiments on this data are instead best interpreted as a domain shift spot check. These annotations are made publicly available in the project’s repository at https://github.com/MikelWL/ICAR.
The 3,000 annotated images are split into 2,100/450/450 (train/val/test) sets. For comparison, we use ICNet [feng2022ic9600] as the baseline; the most established complexity assessment method with viable inference speed for practical deployment among existing approaches. Given this is a binary ‘complex’ or ‘simple’ classification task, we evaluate ConvNeXt-IC in its classification head setting. Table 2 shows the comparative performance.
@@ 3.3 Domain Shift Spot Check on LAION Subset
| ICNet | 0.757 | 0.741 | 0.749 | 0.845 | 397 |
| ConvNeXt-IC | 0.869 | 0.862 | 0.865 | 0.944 | 1744 |
ConvNeXt-IC achieves 0.865 F1-score and 0.944 ROC AUC, outperforming ICNet (0.749 F1, 0.845 ROC AUC) with its 4.4 efficiency advantage (1,744 vs 397 images/second). These results support ConvNeXt-IC’s effective generalisation from curated benchmarks to real-world web distributions, supporting our approach for ICAR’s deployment requirements.
@@ 4.1 Model Design and Dual-Path Training
ICAR integrates ConvNeXt-IC complexity routing with adaptive ViT-L/14 processing. Figure 4 shows the architecture: simple images exit early at layers 8, 12, 16, or 20, while complex images process through all 24 layers. The approach applies dual-path training to produce compatible embeddings from both paths in the same semantic space, addressing the cross-modal alignment challenge in vision-language retrieval.
Architecture: ICAR consists of (1) ConvNeXt-IC for routing decisions (2.45 GFLOPs overhead), (2) ViT-L/14 with early exit points, and (3) identical projection heads ensuring embedding compatibility. Early exit at layer 8 saves 44% computation (108 GFLOPs), while later exits provide efficiency-quality trade-offs. Text processing remains unchanged at 13.3 GFLOPs.

### v2

@@ Abstract
Vision transformers in vision-language models typically use the same amount of compute for every image, regardless of whether it is simple or complex. We propose ICAR (Image Complexity-Aware Retrieval), an adaptive computation approach that enables vision transformers to use less compute for simple images whilst processing complex images through their full network depth. The key challenge is maintaining cross-modal alignment: embeddings from different processing depths must remain compatible for text matching. ICAR solves this through dual-path training that produces compatible embeddings from both the early-exit and full-depth paths. This maintains compatibility between image representations and text embeddings in the same semantic space, whether an image exits early or processes fully. Unlike existing two-stage approaches that require expensive reranking, ICAR enables direct image-text matching without additional overhead. To determine how much compute to use, we develop ConvNeXt-IC, which treats image complexity assessment as a classification task. By applying modern classifier backbones rather than specialised architectures, ConvNeXt-IC achieves state-of-the-art performance, attaining a Pearson correlation coefficient of 0.959 with human labelling whilst delivering 4.4 faster complexity prediction. Evaluated on standard benchmarks augmented with real-world web data, ICAR achieves 20% faster image encoding while maintaining category-level performance and 95% of instance-level performance, enabling sustainable scaling of vision-language systems.
@@ 1 Introduction
Many vision–language retrieval systems use Vision Transformers (ViT) to align images with text queries, enabling applications from web image search [27] to e-commerce product matching [44]. ViTs process images through multiple layers in sequence, progressively refining their representations [28].
@@
Standard ViT systems use a fixed-depth transformer stack and process all image patches uniformly, regardless of image complexity. For example, a 24-layer ViT-L/14 uses the same 175.33 GFLOPs [16] (giga floating-point operations) of compute whether it’s processing a simple product photo or a complex street scene. At web scale, where major platforms handle billions of photos and videos daily [12], this uniform processing can lead to substantial waste of computational resources. Adaptive computation can reduce this inefficiency by allowing simpler images to stop after fewer transformer layers, whilst complex images continue through the full transformer stack.
@@
Adaptive computation for vision-language models faces a fundamental technical challenge in maintaining cross-modal alignment when embeddings are produced from different processing depths. While adaptive computation has been extensively studied for single-modality tasks [37], the adaptive vision-language field remains underexplored. Specifically, no existing work addresses adaptive computation for single-stage image–text retrieval where embeddings from different vision depths remain directly comparable to the text embedding.
@@
This paper proposes Image Complexity-Aware Retrieval (ICAR), which uses image complexity to determine whether an image is routed through fewer transformer layers (early exit) or the full transformer stack (Figure 1). ICAR introduces dual-path training that produces compatible embeddings from both early reduced compute and full processing paths. This enables simple images to use fewer layers whilst complex images process through the full network, all while maintaining cross-modal alignment through unified embedding spaces, eliminating the reranking overhead that existing approaches have.
@@
To support these routing decisions, we develop ConvNeXt-IC (ConvNeXt-Image Complexity). While existing methods treat complexity assessment as a representation learning problem requiring specialised architectures, we instead recognise it as a classification task. We evaluate on both established academic benchmarks (IC9600) and real-world web data (LAION), demonstrating that our approach generalises beyond controlled settings. The following are the contributions of the paper:
@@
Adaptive computation for single-stage image-text retrieval: We introduce ICAR, which maintains cross-modal embedding compatibility across vision depths, enabling direct text matching without reranking overhead.
@@
Reconceptualising complexity assessment as classification: We show that image complexity assessment is better suited as a classification task than representation learning, using modern backbones instead of specialised complexity assessment architectures.
Efficiency–quality trade-off in retrieval: We show that early exits can reduce image encoding compute with minimal impact on retrieval quality, enabled by dual-path training.
@@
This paper is organised as follows: Section 2 discusses related work in image complexity assessment and adaptive computation. Section 3 presents our ConvNeXt-IC complexity detection method with benchmark evaluation on established benchmarks and real-world data. Section 4 details the ICAR architecture and dual-path training approach. Section 5 presents experimental results and analysis of retrieval performance and efficiency gains. Finally, Section 6 concludes with key findings and future directions.
@@ 2.1 Image Complexity Assessment
Image complexity assessment determines how difficult an image is for humans to understand or describe [33]. Psychology research identified key factors: object count, spatial relationships, visual diversity, and background richness [25, 10, 26]. Early computational approaches used hand-crafted features like entropy and compression ratios [34, 23, 31], but these worked only on small datasets.
Current deep learning methods treat complexity assessment as a representation learning problem. ICNet [9] uses a dual-branch architecture combining detail and context pathways with spatial attention modules. ICCORN [13] applies ordinal regression to ICNet features, explicitly modelling the ordinal structure of complexity ratings. MICM [19] introduces motion-inspired assessment using image-to-video generation with hierarchical alignment losses. CLICv2 [21] uses self-supervised pretraining on complexity-related visual patterns.
Related image quality assessment methods explore similar perceptual challenges but focus on quality rather than complexity. HyperIQA [35] employs content-adaptive hyper networks for quality assessment, while CLIPIQA [38] leverages vision-language representations from CLIP [27]. TOPIQ [4] adopts a top-down framework modelling human visual perception characteristics.
These methods follow a common pattern: start with classification baselines, then add complexity-specific architectures for incremental improvements. However, they test only against outdated networks like AlexNet [17] (2012) and ResNet [15] (2016), leaving a large gap unexplored. This raises an open question whether architectural specialisation provides genuine benefits, or whether stronger classification backbones would close the performance gap without the need for additional components. We address this gap using ConvNeXt-V2 (2023) to test whether modern classification architectures outperform specialised complexity assessment designs (Section 3).
@@ 2.2 Adaptive Computation in Machine Learning
Single-modality methods have explored adaptive computation by adjusting processing based on input difficulty. BranchyNet [37] introduced early exits with classifiers at intermediate layers, enabling confident predictions to exit early while difficult cases continue through all layers. Several methods adapt early exits for Vision Transformers through patch processing, token removal, and routing strategies [41, 6, 29, 2, 30].
@@
Limited work exists on adaptive computation for vision-language models. DeeCap [8], MuE [36], and FREE [1] focus on generative settings, while LGViT [40] studies early exiting for vision-only classification. For retrieval, HiVLP [5] uses hierarchical features but requires re-ranking due to separate embedding spaces. In contrast, ICAR targets single-stage image–text retrieval and enforces embedding compatibility across vision depths so that all image embeddings remain directly comparable to the text embedding (Section 4).
@@ 3.1 Proposed Method: ConvNeXt-IC
ConvNeXt-IC predicts whether an image is simple or complex, providing the routing signal for ICAR. Figure 3 shows representative examples from LAION-5B [32].
These examples illustrate the range of complexity that ViTs must handle. Existing image complexity assessment methods treat complexity assessment as a representation learning problem, building specialised architectures to capture complexity patterns [9, 13, 19]. However, the core challenge is discriminating between different levels of visual complexity. We therefore treat complexity assessment as a classification problem, using efficient classification backbones instead of specialised complexity representation architectures.
We therefore fine-tune ConvNeXt-V2-N (224x224) [39], the smallest of the ConvNeXt-V2 ImageNet-22K [7] family of foundation classification models. This choice prioritises efficiency for real-time routing decisions while maintaining the discriminative power needed for accurate complexity assessment. We implement ConvNeXt-IC with two configurations: (1) as a regression model (using a single linear layer predicting continuous scores); and (2) as a binary classification model (using a lightweight two-layer MLP with cross-entropy loss). Both configurations are fine-tuned end-to-end using standard data augmentation (random crops, horizontal flips, colour jittering) on the benchmark datasets.
3.2 Benchmark Evaluation on IC9600
We benchmark ConvNeXt-IC on IC9600 [9] using its regression head, a standard dataset for image complexity assessment. IC9600 contains 9,600 images labelled with complexity scores from 1 (simple) to 5 (complex). Performance is measured following the IC9600 evaluation protocol [9] using standard regression metrics: Pearson Correlation Coefficient (PCC) and Spearman Rank Correlation Coefficient (SRCC) assess how well predicted scores correlate with human ratings, while Root Mean Square Error (RMSE) measures prediction accuracy. Images per second (img/s) inference speed is also reported to assess computational efficiency. Table 1 compares results to existing methods.
@@ 3.2 Benchmark Evaluation on IC9600
| HyperIQA [35] | 0.935 | 0.926 | 0.204 | 61 |
| CLIPIQA [38] | 0.898 | 0.897 | 0.078 | 83 |
| TOPIQ [4] | 0.944 | 0.938 | 0.049 | 125 |
| CLICv2 [21] | 0.933 | 0.927 | - | - |
| ICNet [9] | 0.949 | 0.945 | 0.053 | 397 |
| MICM [19] | 0.953 | 0.943 | 0.049 | 0.01 |
| ICCORN [13] | 0.955 | 0.951 | - | - |
| ConvNeXt-IC | 0.959 | 0.956 | 0.045 | 1744 |
ConvNeXt-IC achieves state-of-the-art performance: 0.959 PCC and 0.956 SRCC, outperforming the previous best method ICCORN (0.955 PCC, 0.951 SRCC) while delivering 1,744 img/s throughput. Most notably, ConvNeXt-IC surpasses methods that were specifically architecturally designed for image complexity detection, like ICNet (0.949 PCC) and MICM (0.953 PCC). This demonstrates that strong, correctly chosen backbones which are then fine-tuned can outperform specialised architectures, providing another instance of the current trend in machine learning where data and compute together can outperform sophisticated architectural design [14, 43].
3.3 Evaluation on a LAION Subset
While IC9600 provides a controlled benchmark dataset, its images are sourced from curated academic datasets. To complement this benchmark, we train and evaluate ConvNeXt-IC on a randomly sampled LAION subset that we manually annotate as ‘complex’ or ‘simple’, reflecting the web distribution where ICAR makes routing decisions. IC9600 serves as the main image complexity prediction evaluation, and this LAION subset complements it by evaluating performance on the web distribution ConvNeXt-IC is intended for.
We sample 3,000 images from LAION-COCO (a subset of LAION-5B) [32] and manually annotate them into complexity classes of ‘complex’ and ‘simple’. The annotation process naturally resulted in the classes being distributed as 49.5% simple and 50.5% complex. Note that, this annotation process was conducted solely by the lead researcher of this paper. Therefore, while ConvNeXt-IC is benchmarked on IC9600 (Section 3.2), results on this LAION subset are best interpreted as an evaluation of performance on the web distribution used for routing.
The 3,000 annotated images are split into 2,100/450/450 (train/val/test) sets using stratified splitting under a fixed random seed (42) for reproducibility. ConvNeXt-IC and the ICNet baseline are trained and evaluated under this split, and results are reported on the held-out test set. For both methods, the prediction threshold is empirically chosen on the validation set and applied to the held-out test set. Table 2 shows the comparative performance.
@@ 3.3 Evaluation on a LAION Subset
| ICNet [9] | 0.864 | 0.823 | 0.843 | 0.924 | 397 |
| ConvNeXt-IC | 0.872 | 0.853 | 0.863 | 0.944 | 1744 |
ConvNeXt-IC achieves 0.863 F1-score and 0.944 ROC AUC, outperforming ICNet (0.843 F1, 0.924 ROC AUC) with its 4.4 efficiency advantage (1,744 vs 397 images/second). These results support the use of ConvNeXt-IC as ICAR’s routing module for web images.
@@ 4.1 Model Design and Dual-Path Training
ICAR integrates an Image Complexity Classifier (ICC), powered by ConvNeXt-IC, with adaptive ViT-L/14 processing. Figure 4 shows the architecture: simple images exit early at layers 8, 12, 16, or 20, while complex images process through all 24 layers. The approach applies dual-path training to produce compatible embeddings from both paths in the same semantic space, addressing the cross-modal alignment challenge in vision-language retrieval.
Architecture: ICAR consists of (1) ICC for routing decisions (powered by ConvNeXt-IC; 2.45 GFLOPs overhead), (2) ViT-L/14 with early exit points, and (3) identical projection heads ensuring embedding compatibility. Early exit at layer 8 saves 44% computation (108 GFLOPs), while later exits provide efficiency-quality trade-offs. Text processing remains unchanged at 13.3 GFLOPs.